# Preparation of Papers for IEEE Sponsored Conferences & Symposia*

### v1

@@ II-C2 The implemented branching rule
Therefore, the implemented branching rule uses the constraint pair $\langle\langle i,a^{i},[t^{i},t^{i}_{u})\rangle,\langle j,a^{j},[t^{j},t^{j}_{u})\rangle\rangle$ for move-move conflicts and $\langle\langle i,a^{i},[t^{i},t^{i}_{u})\rangle,\langle j,v(w^{j}),I^{c}\rangle\rangle$ for move-wait conflicts. No proofs of soundness and solution completeness are provided for CCBS under the implemented branching rule.
II-D Revisiting CCBS
The work in [28] calls the algorithm properties of CCBS into question. The claims are two-fold:
@@ II-D Revisiting CCBS
    under the theoretical branching rule, CCBS does not guarantee termination;
@@
    under the implemented branching rule, CCBS violates property (2).
At its core, the identified issue lies in how wait actions are handled when resolving conflicts.
@@
For Claim 1[28] highlights that under the theoretical branching rule on a wait action $\langle w^{i},t^{i}\rangle$ (where the motion constraint $\langle i,w^{i},[t^{i},t^{i}_{u})\rangle$ forbidding the execution of $w^{i}$ starting at any time in $t^{i},t^{i}_{u})$ is created), only one wait duration $w^{i}_{D}$ out of infinitely many real-valued durations is removed from the search. E.g., a wait action $w=(w^{i}_{\varphi},w^{i}_{D}+\epsilon)$ with $\epsilon$ arbitrarily small is still permitted during $t^{i},t^{i}_{u})$. Therefore, despite satisfying property (2) such that an optimal solution is reachable from the CT root, there may exist infinitely many nodes between them. Hence, CCBS under the theoretical branching rule does not always terminate after finite iterations, despite the existence of a solution. Therefore, CCBS under the theoretical branching rule is not solution complete. In other words, [28] identifies that properties (2) and (3) are not sufficient to guarantee solution completeness.
@@
For Claim 2, [28] provides two counter-examples showing how the implemented branching rule on a move-wait conflict removes valid solutions from the search (violating property (2)), and consequently returning sub-optimal solutions. Thus, CCBS under the implemented branching rule does not guarantee to only return optimal solutions and is therefore not sound.
@@
Furthermore, [28] introduces shifting constraints. They show how a branching rule using shifting constraints, which relies on a parameter $\delta\geq 0$, does not remove collision-free solutions from the search and is therefore an improvement on the implemented branching rule. However, they conclude that if $\delta>0$ then there will always remain joint plans with conflicts, leading to potentially infinitely many nodes being searched. If instead $\delta=0$ then a similar issue as for the theoretical branching rule arises, also potentially requiring the expansion of infinitely many nodes. That is, CCBS using shifting constraints has the same properties as using the theoretical branching rule: sound but without termination guarantees.
III Preliminaries
@@
This section establishes the theoretical foundations for analyzing branching rules in CCBS. We first introduce a decomposed form of CCBS constraints and extend the notion of sound constraint pairs to constraint sets. We then define a sound branching rule as one that preserves all solutions during node expansion and prove that a branching rule is sound if and only if its underlying constraint pair is sound. Using these definitions, we show that a sound branching rule is sufficient to ensure that CCBS is sound, and is also solution complete if CCBS is guaranteed to terminate. Finally, we introduce the intersection interval which captures the complete temporal overlap between a moving agent and a stationary one — useful for designing sound branching rules that handle move-wait conflicts.
III-A Sound Branching
@@
A definition for a sound pair of constraints $\langle c_{1},c_{2}\rangle$ is provided in [20, 31]. Let the set $\mathbf{}{S}$ contain all solutions to a given MAPFR problem, and let $c(\Pi)$ be true if a joint plan $\Pi$ satisfies a constraint $c$, false otherwise.
Definition III.1 (Sound pair of constraints [20, 31]).
@@
For a given MAPFR problem with solutions $\mathbf{}{S}$, a pair of constraints $\langle c_{1},c_{2}\rangle$ is sound if
$c_{1}(\Pi)\vee c_{2}(\Pi),\quad\forall\Pi\in\mathbf{}{S}.$ (4)
@@
Throughout this article, we shall consider a decomposed form of the motion and vertex constraints by describing them as constraint sets comprised of forbidding constraints. These various constraint forms are equivalent, however, the decomposed form allows us to analyze constraint pairs used in branching rules more explicitly than when using motion and vertex constraints. A forbidding constraint $\langle i,a,t\rangle$ forbids $i$ from executing the specific timed action $\langle a,t\rangle$. A motion constraint $c=\langle i,a^{i},[t^{i},t^{i}_{u})\rangle$, which forbids $i$ from executing action $a_{i}$ starting at any time $t\in[t^{i},t^{i}_{u})$, is equivalent to the dense constraint set
@@
such that $c(\Pi)\Leftrightarrow C(\Pi)$ where $C(\Pi)=\bigwedge_{c^{\prime}\in C}c^{\prime}(\Pi)$. We denote this equivalence by $c\Leftrightarrow C$. Similarly, a vertex constraint $c=\langle i,v,I^{c}\rangle$ which forbids $i$ from occupying $v$ at any time $t\in I^{c}$ is equivalent to $C=\left\{\langle i,w,t\rangle\mid v(w)=v,[t,t+w_{D}]\cap I^{c}\neq\varnothing\right\}$. That is, $c\Leftrightarrow C$.
We now provide a corresponding soundness definition for a pair of constraint sets, and then define soundness of a branching rule. Then, in Lemma III.1, we show that a branching rule using a pair of constraint sets is sound if and only if the constraint set pair is sound.
Definition III.2 (Sound pair of constraint sets).
@@
This definition follows quite naturally from extrapolating Definition III.1 to sets of constraints: given a pair of constraints $\langle c_{1},c_{2}\rangle$ and their equivalent constraint sets $C_{1}\Leftrightarrow c_{1}$ and $C_{2}\Leftrightarrow c_{2}$, it holds for all solutions $\Pi\in\mathbf{}{S}$ that
$\quad c_{1}(\Pi)\vee c_{2}(\Pi)$
$\Leftrightarrow$ $\quad C_{1}(\Pi)\vee C_{2}(\Pi)$
$\Leftrightarrow$ $\quad\left(\bigwedge_{c_{1}^{\prime}\in C_{1}}c_{1}^{\prime}(\Pi)\right)\vee\left(\bigwedge_{c_{2}^{\prime}\in C_{2}}c_{2}^{\prime}(\Pi)\right)$
$\Leftrightarrow$ $\quad\forall\langle c_{1}^{\prime},c_{2}^{\prime}\rangle\in C_{1}\times C_{2}:c_{1}^{\prime}(\Pi)\vee c_{2}^{\prime}(\Pi)$
$\Leftrightarrow$ $\quad\forall\langle c_{1}^{\prime},c_{2}^{\prime}\rangle\in C_{1}\times C_{2}:\langle c_{1}^{\prime},c_{2}^{\prime}\rangle\text{ is sound}.$
Definition III.3 (Sound branching rule).
@@
Lemma III.1 (Equivalence of sound branching and sound constraint set pair).
@@
$\mathcal{S}(N^{i})$ $=\left\{\Pi\in\mathbf{}{S}\mid c(\Pi),c\in N^{i}_{C}\right\}$
$=\left\{\Pi\in\mathbf{}{S}\mid c(\Pi),c\in N_{C}\cup\left\{C_{i}\right\}\right\}$
@@
Therefore, for the branching rule to be sound it must hold for every solution $\Pi\in\mathcal{S}(N)$ that $C_{1}(\Pi)\vee C_{2}(\Pi)$. For every $\Pi\in\mathcal{S}(N)$, consider two cases:
@@
For the latter case, we start by assuming that $\langle C_{1},C_{2}\rangle$ is sound. By Definition III.2, every pair $\langle c_{1},c_{2}\rangle\in C_{1}\times C_{2}$ is sound, meaning $c_{1}(\Pi)\vee c_{2}(\Pi)$ holds. Since we assumed $\neg c_{1}(\Pi)$, it follows that $c_{2}(\Pi)$ must be true. This reasoning applies to every pair $\langle c_{1},c_{2}\rangle\in C_{1}\times C_{2}$, such that $\forall c_{2}\in C_{2}:c_{2}(\Pi)$ holds. This means that $C_{2}(\Pi)$ is true and therefore that $\Pi\in\mathcal{S}(N^{2})$. Therefore, if $\langle C_{1},C_{2}\rangle$ is sound then it holds for all $\Pi\in\mathcal{S}(N)$ that $\Pi\in\mathcal{S}(N^{1})\cup\mathcal{S}(N^{2})$. Thus, if $\langle C_{1},C_{2}\rangle$ is sound then the branching rule is sound.
Assume instead that $\langle C_{1},C_{2}\rangle$ is unsound: then by Definition III.2 there exists an unsound pair $\langle c_{1},c_{2}\rangle\in C_{1}\times C_{2}$. Since $\langle c_{1},c_{2}\rangle$ is unsound, there exists a valid solution $\Pi$ such that $\neg(c_{1}(\Pi)\vee c_{2}(\Pi))$, implying $\neg C_{1}(\Pi)$ and $\neg C_{2}(\Pi)$. Therefore, $\Pi\not\in\mathcal{S}(N^{1})\cup\mathcal{S}(N^{2})$. Since $\mathcal{S}(N)$ can contain such a solution $\Pi$, it does not always hold for all $\Pi\in\mathcal{S}(N)$ that $\Pi\in\mathcal{S}(N^{1})\cup\mathcal{S}(N^{2})$. Thus, if $\langle C_{1},C_{2}\rangle$ is unsound then the branching rule is unsound.
Finally, we put branching rules back into the context of the CCBS algorithm by showing in Theorem III.2 that CCBS is sound when using a sound branching rule. Additionally, we show in Theorem III.3 that CCBS is solution complete if it is both sound and guaranteed to terminate.
Theorem III.2 (Soundness).
@@
Theorem III.3 (Solution completeness).
@@
CCBS’s soundness ensures that all solutions are reachable from the CT root. CCBS has exactly one termination condition: when it selects a CT node $N$ such that $N_{\Pi}$ is a solution, $N_{\Pi}$ is returned. CCBS continues expanding nodes until this condition is met. Since all solutions are reachable from the CT root where CCBS begins its search, and CCBS is guaranteed to terminate, it must eventually encounter and return a solution. Therefore, CCBS is solution complete if it is sound and guaranteed to terminate. ∎
Note that Theorem III.2 shows that CCBS is sound if the branching rule is sound. However, it does not show that CCBS is unsound if the branching rule is unsound. Thus, the branching rule being sound is a sufficient but not necessary condition for CCBS to be sound.
III-B Intersection Interval
We now define the intersection interval which will be useful in Section V when introducing a our sound branching rule. Informally, an intersection interval is the time interval when an agent executing a timed move action is in collision with a stationary agent waiting indefinitely at a vertex.
Let a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with collision interval $I^{c}$ be given, illustrated in Figure 4. Suppose now that agent $j$ waits at $v(w^{j})$ for all time instead of only during $t^{j},t^{j}+w_{D}]$. Let all maximally connected time intervals when the agents are in collision be collected in the set of intersection intervals $\bar{\mathbb{I}}$. Since we know that the agents collide, during $^{c}, 𝕀¯\bar{\mathbb{I}}$ contains at least one interval super set of $I^{c}$.
Definition III.4 (Intersection interval).
@@ III-B Intersection Interval
Figure 4 illustrates the relation between $I^{c}$, $\bar{I^{c}}$ and $\bar{\mathbb{I}}$. It is useful to note that $I^{c}$ is the intersection between $j$’s occupancy of vertex $v(w^{j})$ and the intersection interval $\bar{I^{c}}$,
@@
IV Analyzing CCBS’s implemented branching rule
We use the framework from Section III to show that the implemented branching rule is not sound. Therefore, CCBS under the implemented branching rule does not satisfy the sufficient but not necessary condition for soundness from Theorem IV.2: a sound branching rule. Additionally, since the proofs in [20] for CCBS soundness rely on property (2) being satisfied, which is not fulfilled under an unsound branching rule, those proofs are not applicable either. Thus, this section provides complementary evidence to the counter-examples in [28] and Section VI showing that CCBS using the implemented branching rule is indeed not sound.
Lemma IV.1 provides intermediate results that are used in Theorem IV.2 to show that the implemented branching rule is not sound.
Lemma IV.1 (Collision timing for time-shifted move actions).
Let no collision occur when agent $i$ executes $\langle m^{i},t^{i}\rangle$ and $j$ executes $\langle w^{j},t^{j}\rangle$. If $i$ instead executes $\langle m^{i},t^{i}-\epsilon\rangle$ for some $\epsilon>0$, then a collision between $i$ and $j$ executing $\langle m^{i},t^{i}-\epsilon\rangle$ and $\langle w^{j},t^{j}\rangle$ can only occur within the time interval $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$.

### v2

@@ 2.3.2 The Implemented Branching Rule
Therefore, the IBR uses the constraint pair $\langle\langle i,m^{i},[t^{i},t^{i}_{u})\rangle,\langle j,m^{j},[t^{j},t^{j}_{u})\rangle\rangle$ for move-move conflicts and $\langle\langle i,m^{i},[t^{i},t^{i}_{u})\rangle,\langle j,v(w^{j}),\bar{I^{c}}\rangle\rangle$ for move-wait conflicts. No proofs of soundness and solution completeness are provided for CCBS under this branching rule.
2.4 A Critique on CCBS
The work in [29] questions the algorithm properties of CCBS. The claims are two-fold:
@@ 2.4 A Critique on CCBS
    Under the TBR, CCBS does not guarantee termination;
@@
    Under the IBR, CCBS violates property (2).
@@
At its core, the identified issue lies in how wait actions are handled when resolving conflicts. For Claim 1[29] highlights that under the TBR on a wait action $\langle w^{i},t^{i}\rangle$ (giving rise to a motion constraint $\langle i,w^{i},[t^{i},t^{i}_{u})\rangle$), only one wait duration $w^{i}_{D}$ out of infinitely many real-valued durations is removed from the search. E.g., a wait action $w=(w^{i}_{\varphi},w^{i}_{D}+\epsilon)$ with $\epsilon$ arbitrarily small is still permitted during $t^{i},t^{i}_{u})$. Therefore, despite satisfying property (2) ensuring that any optimal solution is reachable from the CT root, there may exist infinitely many nodes between such a solution and the root. Hence, CCBS under the TBR does not always terminate after finite iterations, despite the existence of a solution. Therefore, CCBS under the TBR is not solution complete. Essentially, [29] identifies that properties (2) and (3) are not sufficient to guarantee solution completeness.
@@
For Claim 2, [29] provides two counterexamples showing how the IBR on a move-wait conflict removes valid solutions from the search (violating property (2)), and consequently returning sub-optimal solutions. Thus, CCBS under the IBR does not guarantee to only return optimal solutions and is therefore not sound.
@@
Furthermore, [29] introduces shifting constraints. They show how a branching rule using shifting constraints, which relies on a parameter $\delta\geq 0$, does not remove collision-free solutions from the search and is therefore an improvement on the IBR. However, they conclude that if $\delta>0$ then in the CT there will always remain permitted joint plans containing conflicts, leading to potentially infinitely many nodes being searched. If instead $\delta=0$, then a similar issue as with the TBR arises, also potentially requiring the expansion of infinitely many nodes. That is, CCBS using shifting constraints has the same properties as using the TBR: sound but without termination guarantees.
3 Preliminaries
@@
This section establishes the theoretical foundations for analyzing branching rules in CCBS. We first introduce a decomposed form of CCBS constraints and extend to this the notion of sound constraint pairs. We then define a sound branching rule as one that preserves all solutions during node expansion and prove that a branching rule is sound if and only if its underlying constraint pair is sound. Using these definitions, we show that a sound branching rule is sufficient to ensure that CCBS is sound, and is also solution complete if CCBS is guaranteed to terminate. Finally, we formally introduce the collision interval and intersection interval.
3.1 Sound Branching
@@
A definition for a sound pair of constraints $\langle c_{1},c_{2}\rangle$ is provided in [21, 32]. Let the set $\mathcal{S}$ contain all solutions to a given MAPFR problem, and let $c(\Pi)$ be true if a joint plan $\Pi$ satisfies a constraint $c$, false otherwise.
Definition 3.1 (Sound pair of constraints [21, 32]).
@@
For a given MAPFR problem with solutions $\mathcal{S}$, a pair of constraints $\langle c_{1},c_{2}\rangle$ is sound if
$c_{1}(\Pi)\vee c_{2}(\Pi),\quad\forall\Pi\in\mathcal{S}.$ (4)
@@
Throughout this article, we shall consider a decomposed form of the motion and vertex constraints by describing them as constraint sets comprised of forbidding constraints. These constraint forms are equivalent, however, the decomposed form allows us to analyze constraint pairs used in branching rules more explicitly than when using motion and vertex constraints. A forbidding constraint $\langle i,a,t\rangle$ forbids $i$ from executing the specific timed action $\langle a,t\rangle$. A motion constraint $c=\langle i,a^{i},[t^{i},t^{i}_{u})\rangle$, which forbids $i$ from executing action $a_{i}$ starting at any time $t\in[t^{i},t^{i}_{u})$, is equivalent to the dense constraint set
@@
such that $c(\Pi)\Leftrightarrow C(\Pi)$ where $C(\Pi)=\bigwedge_{c^{\prime}\in C}c^{\prime}(\Pi)$. We denote this equivalence by $c\Leftrightarrow C$. Similarly, a vertex constraint $c=\langle i,v,\bar{I^{c}}\rangle$ forbidding $i$ from occupying $v$ at any time $t\in\bar{I^{c}}$ is equivalent to $C=\left\{\langle i,w,t\rangle\mid v(w)=v,[t,t+w_{D}]\cap\bar{I^{c}}\neq\varnothing\right\}$.
We now provide a corresponding soundness definition for a pair of constraint sets, and then define soundness of a branching rule. Then, in Lemma 3.1, we show that a branching rule using a pair of constraint sets is sound if and only if the constraint set pair is sound.
Definition 3.2 (Sound pair of constraint sets).
@@
Definition 3.3 (Sound branching rule).
@@
Lemma 3.1 (Equivalence of sound branching and sound constraint set pair).
@@
$\mathcal{S}(N^{i})$ $=\left\{\Pi\in\mathcal{S}\mid c(\Pi),c\in N^{i}_{C}\right\}$
$=\left\{\Pi\in\mathcal{S}\mid c(\Pi),c\in N_{C}\cup\left\{C_{i}\right\}\right\}$
@@
Therefore, for the branching rule to be sound it must hold for every solution $\Pi\in\mathcal{S}(N)$ that $C_{1}(\Pi)\vee C_{2}(\Pi)$ since then $\mathcal{S}(N)=\mathcal{S}(N^{1})\cup\mathcal{S}(N^{2})$. For every $\Pi\in\mathcal{S}(N)$, consider two cases:
@@
For the latter case, we start by assuming that $\langle C_{1},C_{2}\rangle$ is sound. By Definition 3.2, every pair $\langle c_{1},c_{2}\rangle\in C_{1}\times C_{2}$ is sound, meaning $c_{1}(\Pi)\vee c_{2}(\Pi)$ holds. Since we assumed $\neg c_{1}(\Pi)$, it follows that $c_{2}(\Pi)$ must be true. This reasoning applies to every pair $\langle c_{1},c_{2}\rangle\in C_{1}\times C_{2}$, such that $\forall c_{2}\in C_{2}:c_{2}(\Pi)$ holds. This means that $C_{2}(\Pi)$ is true and therefore that $\Pi\in\mathcal{S}(N^{2})$. Therefore, if $\langle C_{1},C_{2}\rangle$ is sound then it holds for all $\Pi\in\mathcal{S}(N)$ that $\Pi\in\mathcal{S}(N^{1})\cup\mathcal{S}(N^{2})$. Thus, if $\langle C_{1},C_{2}\rangle$ is sound then the branching rule is sound.
Assume instead that $\langle C_{1},C_{2}\rangle$ is unsound, then by Definition 3.2 there exists an unsound pair $\langle c_{1},c_{2}\rangle\in C_{1}\times C_{2}$. Since $\langle c_{1},c_{2}\rangle$ is unsound, there exists a valid solution $\Pi$ such that $\neg(c_{1}(\Pi)\vee c_{2}(\Pi))$, implying $\neg C_{1}(\Pi)$ and $\neg C_{2}(\Pi)$. Therefore, $\Pi\not\in\mathcal{S}(N^{1})\cup\mathcal{S}(N^{2})$. Since $\mathcal{S}(N)$ can contain such a solution $\Pi$, it does not always hold for all $\Pi\in\mathcal{S}(N)$ that $\Pi\in\mathcal{S}(N^{1})\cup\mathcal{S}(N^{2})$. Thus, if $\langle C_{1},C_{2}\rangle$ is unsound then the branching rule is unsound.
Finally, we put branching rules back into the context of the CCBS algorithm by showing in Theorem 3.2 that CCBS is sound when using a sound branching rule. Additionally, we show in Theorem 3.3 that CCBS is solution complete if it is both sound and guaranteed to terminate.
Theorem 3.2 (Soundness).
@@
Theorem 3.3 (Solution completeness).
@@
CCBS’s soundness ensures that all solutions are reachable from the CT root. CCBS has exactly one termination condition: when it selects a CT node $N$ such that $N_{\Pi}$ is a solution, then $N_{\Pi}$ is returned. CCBS continues expanding nodes until this condition is met. Since all solutions are reachable from the CT root where CCBS begins its search, and CCBS is guaranteed to terminate, it must eventually encounter and return a solution. Therefore, CCBS is solution complete if it is sound and guaranteed to terminate. ∎
Note that Theorem 3.2 shows that CCBS is sound if the branching rule is sound. However, it does not show that CCBS is unsound if the branching rule is unsound. Thus, the branching rule being sound is a sufficient but not necessary condition for CCBS to be sound.
3.2 Collision Interval and Intersection Interval
We now formally define the collision interval and intersection interval, which will be useful in Section 5 when introducing $\delta$-BR. Given a conflict $\langle\langle a^{i},t^{i}\rangle,\langle a^{j},t^{j}\rangle\rangle$, the collision interval $I^{c}$ is the first contiguous time interval during which agents $i$ and $j$ are in collision with each other while executing their respective timed actions. Such a collision interval must exist for any move-move or move-wait conflict, otherwise it is not a conflict. On the other hand, an intersection interval $\bar{I^{c}}$ exists only for a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$. Informally, for such a move-wait conflict, $\bar{I^{c}}$ is the time interval during which $i$ and $j$ are in collision if $j$ remains stationary at vertex $v(w^{j})$ for all time instead of only during the time when executing $\langle w^{j},t^{j}\rangle$. If multiple such intervals exist, then $\bar{I^{c}}$ is the interval containing $I^{c}$.
More concretely, let a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with collision interval $I^{c}$ be given (illustrated in Figure 4). Suppose now that agent $j$ waits at $v(w^{j})$ for all time instead of only during $t^{j},t^{j}+w_{D}]$. Let all maximally connected time intervals when the agents are in collision be collected in the set of intersection intervals $\bar{\mathbb{I}}$. Since we know that the agents collide, during $^{c}, 𝕀¯\bar{\mathbb{I}}$ contains at least one interval super set of $I^{c}$.
Definition 3.4 (Intersection interval).
@@ 3.2 Collision Interval and Intersection Interval
Figure 4 illustrates the relation between $I^{c}$, $\bar{I^{c}}$ and $\bar{\mathbb{I}}$. It is useful to note that the collision interval $I^{c}$ is the intersection between $j$’s occupancy of vertex $v(w^{j})$ and the intersection interval $\bar{I^{c}}$,
@@
4 Analyzing CCBS’s Implemented Branching Rule
We use the framework from Section 3 to show that the IBR is not sound. Therefore, CCBS using the IBR does not satisfy the sufficient but not necessary condition for soundness from Theorem 4.2: a sound branching rule. Additionally, since the proofs in [21] for CCBS’s soundness rely on property (2) being satisfied, which is not fulfilled under the IBR, those proofs do not apply. Thus, this section provides complementary evidence to the counterexamples in [29] and Section 6 showing that CCBS using the IBR is indeed not sound.
Lemma 4.1 provides intermediate results that are used in Theorem 4.2 to show that the IBR is not sound.
Lemma 4.1 (Collision timing for time-shifted move actions).
Let $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ not be a conflict. If $i$ instead executes $\langle m^{i},t^{i}-\epsilon\rangle$ for some $\epsilon>0$, then a collision with $j$ executing $\langle w^{j},t^{j}\rangle$ can only occur within the time interval $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$.